OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x2 - 24x + 32? Answers: (-2, -8) (-2, 8) (2, -8) (2, 8) My answer: C

hartnn (hartnn):

nopes ..

hartnn (hartnn):

u know how to solve ?

OpenStudy (anonymous):

only imaginary root possible.........

OpenStudy (anonymous):

I'm stuck between finding out if they're both positives?

OpenStudy (anonymous):

how to solve?

hartnn (hartnn):

2(3x^2-12x+16) =2 (3[x^2-4x+4-4]+16) =2(3[(x-2)^2]-12+16) =2(3[(x-2)^2]) +8 so the vertex is (2,8)

OpenStudy (anonymous):

YES!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1 :) Thank you, that's what I thought when I re-did it :)

OpenStudy (anonymous):

ok my mistake....

hartnn (hartnn):

welcome ^_^